OpenStudy (anonymous):

Use elimination to solve the following system of equations. Enter your answer as an ordered triple, like this example: (2, -3, 4). {4x+y+3z=0 2x-2y-z=10 3x - 2y + 2z = 11} please help ! will award a medal !

OpenStudy (anonymous):

( 1, -4, 0 )

OpenStudy (ankitshaw):

subtract eqn 3 and 2, also multiply 2 to eqn 1 and add it to eqn 2... doing this you will get two eqn with variable y eliminated... again use elimination in those two eqn to get your answer.

OpenStudy (anonymous):

(1, -4, O) First add the second and third equation to make an equation with two variables, next subtract the first equation from the second in order to get another equation with two variables. Now you should have a system of tw equations with two variables: -x - 3z = -1 10x - 7z = 10 Now multiply the first equation by 10 on both sides of equal sign add the result to second equation: -37z = 0 z = 0 Plug the value into the 10x - 7z = 10 equation to determine the X-value: 10x - 7(0) = 10 10x = 10 X = 1 Lastly, plug the x- and z-value into any equation in first system in order to determine the y-value: 4(1) + y + 3(0) = 0 4 + y + 0 = 0 4 + y = 0 y = -4 Thus the answer is {1, -4, 0}